OpenStudy (anonymous):

Given the function g(x) = −3x + 4, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. A: The value of g(−2) is smaller than the value of g(4). B: The value of g(−2) is the same as the value of g(4). C: The values of g(−2) and g(4) cannot be compared. D: The value of g(−2) is larger than the value of g(4).

OpenStudy (anonymous):

This is easy, look: It's just asking you to plug in x = -2 and x = 4 into your equation g(x) = −3x + 4: x = -2 g(x) = −3x + 4 g(x) = -3(-2) + 4 g(x) = 6 + 4 g(x) = 10 x = -4 g(x) = −3x + 4 g(x) = -3(4) + 4 g(x) = -12 + 4 g(x) = -8 Can you tell what the correct answer is?

OpenStudy (anonymous):

You now know that g(-2) = 10 and g(4) = -8.

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

No, look again.

OpenStudy (anonymous):

Is 10 less than -8? Because that's what A states.

OpenStudy (anonymous):

No, So its D

OpenStudy (anonymous):

Yes, you got it.

OpenStudy (anonymous):

Thank you